• There were ten red bottles sitting on the wall. The probability of a red bottle accidentally falling is 0.95. What is the probability that fewer than 8 of the green bottles accidentally fall?
1
Expert's answer
2022-06-16T09:51:28-0400

Since there are no green bottles, then  the probability that fewer than 8 of the green bottles accidentally fall is "0."


Let "X=" the number of red bottles which fall: "X\sim Bin (n, p)."

Given "n=10, p=0.95, q=1-p=0.05."



"P(X<8)=1-P(X=8)-P(X=9)""-P(X=10)=1-\dbinom{10}{8}(0.95)^{8}(0.08)^{10-8}""-\dbinom{10}{9}(0.95)^{9}(0.08)^{10-9}-\dbinom{10}{10}(0.95)^{10}(0.08)^{10-10}""=0.0115"